OpenStudy (anonymous):

Find the rectangular coordinates of the point with the polar coordinates (3, 3 divided by 2 pi). (0, 3) (0, -3) (-3, 0) (3, 0)

ganeshie8 (ganeshie8):

\(\large x = r\cos \theta\) \(\large y = r\sin\theta\)

OpenStudy (anonymous):

what happens next?

ganeshie8 (ganeshie8):

plugin the given values

OpenStudy (anonymous):

what is r equal to?

ganeshie8 (ganeshie8):

you're given polar coordinates : \(\large (3, \frac{3 \pi}{2})\)

ganeshie8 (ganeshie8):

that means, you're given the length and angle : \(\large r = 3\), \(\large \theta = \frac{3\pi}{2}\)

ganeshie8 (ganeshie8):

use the given formula to convert it to rectangular form

OpenStudy (anonymous):

ohhh okay so it would be x: cos 3pi/2 and y would be: 3 sin 3pie/2

ganeshie8 (ganeshie8):

yup !

OpenStudy (anonymous):

sorry redo : x = 3 cos 3pi/2 y = 3 sin 3pi/2

ganeshie8 (ganeshie8):

it doesnt matter, but small correction : ohhh okay so it would be x: 3cos 3pi/2 and y would be: 3 sin 3pie/2

ganeshie8 (ganeshie8):

yes...

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

how do u simplify that?

ganeshie8 (ganeshie8):

whats cos (3 pi/2) ?

OpenStudy (anonymous):

.99

ganeshie8 (ganeshie8):

^^ look at that image 3pi/2 takes u to negative y axis

OpenStudy (anonymous):

so it is equal to -1? or 0?

OpenStudy (anonymous):

i think its 0

OpenStudy (anonymous):

btw thank u @myininaya

ganeshie8 (ganeshie8):

yes ! on unit circle, "cos" represents "x coordinate" "sin" represents "y coordinate"

OpenStudy (anonymous):

ok so is the answer (0,3)

ganeshie8 (ganeshie8):